OpenStudy (anonymous):

Factor compeletly. Remember to look first for a common factor and check by multiplying. If a polynomial is prime,state this. -36a^2 -96ab -64b^2

OpenStudy (anonymous):

Common factors first, then work the polynomial.....\[-36a^2 -96ab -64b^2=-4(9a^2+24ab+16b^2)=-4(3a+4b)^2\]It is really important on problems like this one to recognize the patterns from in the book, in particular for this problem\[A^2+2AB+B^2=(A+B)^2\]